OpenStudy (anonymous):

Solve (1/n2) + (1/n) = (1/2n2)

OpenStudy (anonymous):

\[\frac{ 1 }{ n^{2}} + \frac{ 1 }{ n } = \frac{ 1 }{ 2n^{2} } / \times 2n^{2} => 2 + 2n = 1 => 2n = -1 => n = \frac{ -1 }{ 2 }\]